# An effective $\boldsymbol{\Lambda}$-Szekeres modelling of the local Universe with Cosmicflows-4

### v1

@@ 2.4 Quasilocal expansion field of the cosmic neighbourhood
An intrinsic limitation of existing velocity field reconstructions is due to small scale nonlinear motions and gravitational clustering, which can only be partially modelled in the reconstruction process [^2] For example by adding a constant nonlinear component to the velocity dispersion which is treated as a free parameter [Courtois_2023, Valade_2024, Courtois_2025]. unavoidably contaminating the linear reconstruction. As a consequence, for example, there are multiple shells in which the reconstructed linear density contrast is $\delta\geq 6$ (with a probability of occurring for a single cell of order $10^{-4}\%$), or even below the minimum $\delta<-1$. Therefore, to construct a multistructured $\Lambda$-Szekeres effective model of the local Universe that leverages the linear (and quasi-linear [^3] As proven in [GasparBuchert_2021, GasparBuchertOstrowski_2023], the $\Lambda$-Szekeres models are exact solutions of the first-order scheme of the Generalised Relativistic Zel’dovich Approximation (GRZA). Therefore, since the GRZA—similar to its Newtonian counterpart—remains valid beyond the linear regime of the density contrast at first order, we can expect a first-order $\Lambda$-Szekeres approach to also provide a reliable description within the quasi-linear regime. ) regime of the HAMLET CF4 reconstruction, we must further process the data to mitigate nonlinear contamination. To achieve this, we employ a smoothing algorithm that computes averages over groups of neighboring cells for both the density contrast and PV fields, thereby filtering out small-scale nonlinear fluctuations while retaining the large-scale structure relevant to our analysis. Specifically, we apply a coarse-graining procedure over groups of 16 voxels—equivalent to increasing the voxel side length from $\approx 4~\mathrm{Mpc/h}$ to $\approx 64~\mathrm{Mpc/h}$—which reduces the density contrast to the range $0.3\leq\delta\leq 0.7$. This level of smoothing ensures that the reconstructed fields remain within the quasi-linear regime whilst preserving the main morphological features of the CF4 structures. We thus perform the multistructured $\Lambda$-Szekeres fitting over the CF4 coarse-grained reconstruction data following the procedure detail in subsections 2.2 and 2.3.
@@
In Fig. 1, we plot, in the three fundamental planes, the ratio between the reconstructed quasilocal expansion field, $H_{q}$, and the underlying Hubble parameter, $H_{0}$. We find predicted variations in the Hubble field absolute magnitude of up to 10%, with a non-trivial spatial profile ultimately reflecting the anisotropy and inhomogeneity of the local cosmic web. In particular, along the SGX–SGY and SGX–SGZ planes, we observe an essentially dipole-like feature in the expansion field, which is absent in the SGY–SGZ plane, further highlighting the directional dependence of the local Hubble flow within the local Universe. Moreover, we note that this behaviour is consistent with recent independent analyses reporting a dipole-like pattern in the local expansion field inferred from CF4 data [Salzano_2025].
@@ 3 Impact on cosmological inference
We are now in a position to assess how the multi-$\Lambda$-Szekeres description of the cosmic environment affects cosmological observations. As a testbed, we use the subsample of Pantheon+ type Ia supernovae (SNe Ia) in the redshift range $0.023<z<0.15$, i.e., the same range employed by the SH0ES collaboration [Riess_2022] to infer the value of the Hubble constant.
Note that, a priori, the different local expansion rates predicted by the patchwork of multi-structured $\Lambda$-Szekeres models lead to corresponding anisotropic and inhomogeneous modifications in the distance–redshift relation. Here, following the findings of Sec. 2.4, we can approximate the luminosity distance-redshift relationship predicted by our model within the volume of the used CF4 reconstruction by exclusively using the quasilocal averages of the expansion field, yielding (see also App. B)
@@
To quantify the significance of these deviations, we adopt a bootstrapping approach and generate one thousand mock realisations of the CF4 smoothed reconstructions, each of which is then modelled using the proposed multi-$\Lambda$-Szekeres description. For each CF4 realisation, we extract a corresponding $H_{q}$ map, from which we compute the luminosity-distance corrections for the SNe Ia sample. For each supernova, we then assign the correction to the luminosity distance—and its associated uncertainty—as the mean, $\bar{\Delta}d_{L}^{\Lambda\text{-Sk}}$, and standard deviation, $\sigma_{d_{L}}^{\mathrm{Sk}}$, of the corrected distances across all realisations of the HAMLET CF4 reconstructions.
In Fig. 3 we show the comparison of the $d_{L}$ predictions for the SNe Ia sample obtained from the $\Lambda$-Szekeres reconstruction (averaged over all the realisations), a fiducial FLRW cosmology, [^5] Here, the fiducial distances in the FLRW cosmology for a given $z$ are computed using a standard cosmographic expansion up to quadratic order in redshift. We fix the cosmographic parameters to $H_{0}=73~\mathrm{km\,s^{-1}\,Mpc^{-1}}$ and $q_{0}=-0.45$. and a PV-corrected FLRW background. Interestingly, the distribution of luminosity distances in the $\Lambda$-Szekeres description directly suggests an average infall within the Pantheon+ SNe Ia sub-sample at redshift $z\approx 0.04$, as was also noted in [Sorrenti_2024] following a cosmographic analysis of the Pantheon+ sample. Furthermore, we see that the $\Lambda$-Szekeres predictions align with those of the PV-corrected fiducial cosmology at the lowest redshifts, while they increasingly deviate as $z$ grows, signalling the cumulative impact of inhomogeneities on the light propagation. This divergence reflects the fact that quasilocal variations in the expansion field can accumulate with distance—mirroring the anisotropic geometry of the local Universe—and thereby producing progressively stronger departures from the homogeneous FLRW expectation at higher redshifts along specific directions.
@@
We perform the analysis using three covariance matrices: (i) the standard SH0ES covariance without PV corrections; (ii) the SH0ES covariance including PV corrections from the 2M++ velocity–field reconstruction; and (iii) the former augmented by adding in quadrature the standard deviation of the $\Lambda$-Szekeres corrections, $\sigma_{d}^{\mathrm{\Lambda-Sk}}$. The resulting constraints [^6] Here, we monitor the convergence following the prescription of emcee, estimating the autocorrelation time $\tau$ every 100 steps and declaring convergence once $\tau$ changes by less than 1%. are presented in Fig. 4. Consistent with [Riess_2022], including PV corrections reduces the inferred distances and increases the $H_{0}$ best fit from $H_{0}=72.46\begin{subarray}{c}+0.82\\ -0.51\end{subarray}$ to $H_{0}=73.19\pm 0.49$. A qualitatively similar trend is obtained when including the $\Lambda$-Szekeres corrections, yielding $H_{0}=72.9\begin{subarray}{c}+0.81\\ -0.49\end{subarray}$. The latter result, consistent with the findings of [Giani_2024a] regarding the impact of Laniakea on cosmological inference, suggests that the Hubble tension cannot be resolved solely by modelling the expansion rate field within the local Universe. On the contrary, accounting for the local expansion field—even through a fully anisotropic treatment of the cosmic neighbourhood—appears to exacerbate the tension. This points toward the presence of an overall overdensity, rather than an underdensity, within the local cosmic web.
Finally, it is natural to ask whether the effective–model corrections behave as random scatter or exhibit a coherent structure. As shown in Fig. 5, the corrections, clearly display a directional pattern rather than a random distribution. To investigate further, in Fig. 6 we show the expected distance corrections at $z=0.03,\,0.06,\,\text{and},\,0.09$. These corrections—reflecting the geometry of the underlying quasilocal expansion rate field—exhibit a directional pattern, signalling the alternation of cosmic structures, i.e., walls and voids, within the local Universe. As such, they reveal the nontrivial geometry of the local cosmic web and its influence on distance measurements, highlighting the a priori importance of accounting for anisotropic expansion effects in cosmological analyses within the cosmic neighbourhood.
@@
Figure 6: Sky map of the $\Lambda$-Szekeres distance corrections as in Fig.5, but shown in different redshift bins. One can see that, especially at low redshifts, the corrections follow a non-random pattern across the sky.
@@ 4 Conclusions
To constrain our model, we adopt a data–driven strategy based on the HAMLET CF4 reconstruction [Valade_2022, Valade_2024] of the local density contrast and the full three–dimensional peculiar velocity field out to redshift $z\lesssim 0.15$ [Tully_2023]. To suppress spurious small–scale nonlinearities, we smooth the original fields using a coarse–graining procedure over neighbouring voxels, ensuring that the resulting maps remain fully within the quasilinear regime. We then leverage the quasilocal formalism of the $\Lambda$–Szekeres solutions [Sussman_2012] to fit an approximate model to these smoothed data, and extract the quasilocal averages of both the expansion field and the spatial curvature in the local Universe (see Figs. 1 and 2).
@@
To assess the impact of these corrections on cosmological inference, we used the Pantheon+ SNe Ia subsample in the redshift range 0.023<z<0.150.023<z<0.15—the same range employed by the SH0ES collaboration to infer the Hubble constant. We performed a bootstrapping procedure over realisations of the employed CF4 reconstruction—and hence of the extracted quasilocal average expansion field—to determine, for each supernova in the sample, the directional corrections to the luminosity distances relative to the fiducial FLRW cosmology, along with their associated uncertainties. These corrections—which we find can modify supernova distances by up to 10%—were then incorporated into an MCMC analysis of a second-order cosmographic expansion’s parameter space to quantify the bias introduced by the local Universe expansion field on the estimation of $H_{0}$. We found that the inferred value of $H_{0}$ shifts upward when considering the corrections predicted by our approximate $\Lambda$–Szekeres model, worsening rather than relieving the Hubble tension.
These findings align with a previous analysis [Giani_2024a], which suggests that local structures cannot reconcile early- and late-Universe determinations of $H_{0}$. However, while [Giani_2024a] considered only a homogeneous, anisotropic model of the Laniakea supercluster, we account for the full network of filaments, clusters, and voids in the local Universe, providing a more complete assessment of how local structure affects supernova distances and, consequently, the inferred value of $H_{0}$. Nonetheless, we find that including the full complexity of the local cosmic web does not seem to alleviate the Hubble tension, indicating that the solution might lie beyond local-structure effects.
@@ Appendix A The quasilocal formalism for $\Lambda$-Szekeres models
The quasilocal formalism of $\Lambda$-Szekeres models developed Sussman and Bolejko in [Sussman_2012] and expanded upon in [Sussman_2015, Sussman_2016, Sussman_2017], has been demonstrated to be particularly well adapted to describe multi-structured $\Lambda$-Szekeres models [Sussman_2015, Sussman_2016].
Within this formalism it is convenient to write $\Lambda$-Szekeres metric in spherical coordinates as [Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017]
@@
where $a=a(t,r)$, $\Gamma=\Gamma(t,r)=1+ra^{\prime}/a$, $a^{\prime}=\partial_{r}a$, $\mathcal{K}_{q,\text{ini}}=\mathcal{K}_{q,\text{ini}}(r)$ is an a priori arbitrary function of $r$, whilst the Szekeres dipole function $\boldsymbol{W}(r,\theta,\phi)$ and its magnitude $\mathcal{W}(r)$ are given by
@@
where $X,Y$ and $Z$ are arbitrary functions of the radial variable. We can then define for the Hubble scalar, $H$, the curvature scalar, $\mathcal{K}$, and the density scalar, $\rho$, their respective “q-scalars” via [Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017]
@@
Therefore, the q-scalars are proper-volume averages of their corresponding scalars with weight factor $\mathcal{F}$, which can be interpreted as a measure of the total gravitational binding energy within the volume [^7] This can be seen directly for spherically symmetric systems in the Newtonian limit, for which the identification is exact [Hayward_1996].. We can also define the exact fluctuations of any dynamical scalar $A$ as
@@
It can then be shown that the dynamics of the Szekeres solutions is now completely determined by the set of first-order evolution equations [Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017]
@@
where we have defined $\Delta^{(\rho)}:=D^{(\rho)}/\rho_{q}$. The evolution equations (A.9)-(A.12) are completed by the following algebraic constraints on the dynamical variables [Sussman_2012, Sussman_2015, Sussman_2016, Sussman_2017]
@@
Appendix B Code routine validation
To validate our code for the $\Lambda$-Szekeres modelling, we employ a simple simulation of a toy model, the so-called Hubble Bubble, whose analytical behaviour is well understood. Specifically, we consider a spherical region of radius $r_{b}$ with a constant linear density contrast $\delta$, embedded in an otherwise flat FLRW background. The variation of the Hubble parameter inside the bubble is related to the density contrast and to the background expansion rate by [Marra_2013]
$\frac{\delta H}{\bar{H}}=-\frac{\delta}{3}\Theta(\delta,\Omega_{m})f(\Omega_{m})\;,$ (B.1)
where $\bar{H}$ is the Hubble parameter outside the density contrast bubble, $\Theta\approx 1$ in the linear regime, and $f\approx\Omega_{m}^{0.55}$. To validate or modelling framework, we compare the distance–redshift predictions for light sources for (i) the effective $\Lambda$-Szekeres model, (ii) Newtonian perturbation theory, and (iii) the standard FLRW model. However, to build a Hubble diagram and compare the various models predictions, we need to specify the underlying distance–redshift relationship in the Hubble bubble model.

### v2

@@ 2.4 Quasilocal expansion field of the cosmic neighbourhood
An intrinsic limitation of existing velocity field reconstructions is due to small scale nonlinear motions and gravitational clustering, which can only be partially modelled in the reconstruction process [^2] For example by adding a constant nonlinear component to the velocity dispersion which is treated as a free parameter [24, 111, 25]. unavoidably contaminating the linear reconstruction. As a consequence, for example, there are multiple shells in which the reconstructed linear density contrast is $\delta\geq 6$ (with a probability of occurring for a single cell of order $10^{-4}\%$), or even below the minimum $\delta<-1$. Therefore, to construct a multistructured $\Lambda$-Szekeres effective model of the local Universe that leverages the linear (and quasi-linear [^3] As proven in [30, 29], the $\Lambda$-Szekeres models are exact solutions of the first-order scheme of the Generalised Relativistic Zel’dovich Approximation (GRZA). Therefore, since the GRZA — similar to its Newtonian counterpart — remains valid beyond the linear regime of the density contrast at first order, we can expect a first-order $\Lambda$-Szekeres approach to also provide a reliable description within the quasi-linear regime. ) regime of the HAMLET CF4 reconstruction, we must further process the data to mitigate nonlinear contamination. To achieve this, we employ a smoothing algorithm that computes averages over groups of neighboring cells for both the density contrast and PV fields, thereby filtering out small-scale nonlinear fluctuations while retaining the large-scale structure relevant to our analysis. Specifically, we apply a coarse-graining procedure over groups of $16\times 16\times 16$ voxels — equivalent to an effective coarse graining scale of $\simeq 64~\mathrm{Mpc/h}$ — which reduces the density contrast to the range $0.3\leq\delta\leq 0.7$. This level of smoothing ensures that the reconstructed fields remain within the quasi-linear regime whilst preserving the main morphological features of the CF4 structures (see also App. B). We thus perform the multistructured $\Lambda$-Szekeres fitting over the CF4 coarse-grained reconstruction data following the procedure detailed in subsections 2.2 and 2.3.
@@
In Fig. 1, we plot the ratio between the reconstructed quasilocal expansion field, $H_{q}$, and the underlying Hubble parameter, $H_{0}$, in the three fundamental planes. We find predicted variations in the Hubble field absolute magnitude of up to 10%, with a non-trivial spatial profile ultimately reflecting the anisotropy and inhomogeneity of the local cosmic web. In particular, along the SGX–SGY and SGX–SGZ planes, we observe an essentially dipole-like feature in the expansion field, which is absent in the SGY–SGZ plane, further highlighting the directional dependence of the local Hubble flow within the local Universe. Moreover, we note that this behaviour is consistent with recent independent analyses reporting a dipole-like pattern in the local expansion field inferred from CF4 data [92].
@@ 3 Impact on cosmological inference
We are now in a position to assess how the multi-$\Lambda$-Szekeres description of the cosmic environment affects cosmological observations. As a testbed, we use the subsample of Pantheon+ type Ia supernovae (SNe Ia) in the redshift range $0.023<z<0.15$, i.e., the same range employed by the SH0ES collaboration [91] to infer the value of the Hubble constant.
Note that, a priori, the different local expansion rates predicted by the patchwork of multi-structured $\Lambda$-Szekeres models lead to corresponding anisotropic and inhomogeneous modifications in the distance–redshift relation. Here, following the findings of Sec. 2.4, we can approximate the luminosity distance-redshift relationship predicted by our model within the volume of the used CF4 reconstruction by exclusively using the quasilocal averages of the expansion field, yielding (see also App. C)
@@
To quantify the significance of these deviations, we adopt a bootstrapping approach and generate one thousand mock realisations of the CF4 smoothed reconstructions. As the CF4 coarse-grained fields provide a mean value and a $1\sigma$ uncertainty for both the density contrast and PV fields in each voxel, we build an ensemble by sampling these quantities independently voxel-by-voxel (i.e., a Monte–Carlo error propagation at the field level). We then model each CF4 realisation using the proposed multi-$\Lambda$-Szekeres description, and extract a corresponding $H_{q}$ map, from which we compute the luminosity-distance corrections for the SNe Ia sample. For each supernova, we then assign the correction to the luminosity distance and its associated uncertainty as the mean, $\bar{\Delta}d_{L}^{\Lambda\text{-Sk}}$, and standard deviation, $\sigma_{\Delta d_{L}}^{\mathrm{Sk}}$, respectively, of the corrected distances across all realisations of the HAMLET CF4 reconstructions.
In Fig. 3 we show the comparison of the $d_{L}$ predictions for the SNe Ia sample obtained from the $\Lambda$-Szekeres reconstruction (averaged over all the realisations), a fiducial FLRW cosmology [^5] Here, the fiducial distances in the FLRW cosmology for a given $z$ are computed using a standard cosmographic expansion up to second order in redshift. We fix the cosmographic parameters to $H_{0}=73~\mathrm{km\,s^{-1}\,Mpc^{-1}}$ and $q_{0}=-0.45$. , and a PV-corrected FLRW background. Interestingly, the distribution of luminosity distances in the $\Lambda$-Szekeres description directly suggests an average infall within the Pantheon+ SNe Ia sub-sample at redshift $z\approx 0.04$, as was also noted in [99] following a cosmographic analysis of the Pantheon+ sample. Furthermore, we see that the $\Lambda$-Szekeres predictions align with those of the PV-corrected fiducial cosmology at the lowest redshifts, while they increasingly deviate as $z$ grows, signalling the cumulative impact of inhomogeneities on the light propagation. This divergence reflects the fact that quasilocal variations in the expansion field can accumulate with distance—mirroring the anisotropic geometry of the local Universe—and thereby producing progressively stronger departures from the homogeneous FLRW expectation at higher redshifts along specific directions.
@@
We perform the analysis using three covariance matrices: (i) the standard SH0ES covariance without PV corrections; (ii) the SH0ES covariance including PV corrections from the 2M++ velocity–field reconstruction; and (iii) the former augmented by adding in quadrature the standard deviation of the $\Lambda$-Szekeres corrections, $\sigma_{\Delta d_{L}}^{\mathrm{\Lambda-Sk}}$. The resulting constraints [^6] Here, we monitor the convergence following the prescription of emcee, estimating the autocorrelation time $\tau$ every 100 steps and declaring convergence once $\tau$ changes by less than 1%. are presented in Fig. 4. Consistent with [91], including PV corrections reduces the inferred distances and increases the $H_{0}$ best fit from $H_{0}=72.46\begin{subarray}{c}+0.82\\ -0.51\end{subarray}$ to $H_{0}=73.19\pm 0.49$. A qualitatively similar trend is obtained when including the $\Lambda$-Szekeres corrections, yielding $H_{0}=72.9\begin{subarray}{c}+0.81\\ -0.49\end{subarray}$. The latter result, consistent with the findings of [43] regarding the impact of Laniakea on cosmological inference, suggests that the Hubble tension cannot be resolved solely by modelling the expansion rate field within the local Universe. On the contrary, accounting for the local expansion field — even through a fully anisotropic treatment of the cosmic neighbourhood — appears to exacerbate the tension. This points toward the presence of an overall overdensity, rather than an underdensity, within the local cosmic web.
Finally, it is natural to ask whether the effective–model corrections behave as random scatter or exhibit a coherent structure. As shown in Fig. 5, the corrections, clearly display a directional pattern rather than a random distribution. To investigate further, in Fig. 6 we show the expected distance corrections at $z=0.03,\,0.06,\,\text{and},\,0.09$. These corrections, which reflect the geometry of the underlying quasilocal expansion rate field, exhibit a directional pattern that indicates an alternating arrangement of cosmic stuctures, i.e., walls and voids, within the local Universe. As such, they reveal the nontrivial geometry of the local cosmic web and its influence on distance measurements, highlighting the a priori importance of accounting for anisotropic expansion effects in cosmological analyses within the cosmic neighbourhood. This interpretation is further supported by Fig. 7, where we show the dispersion $\sigma_{\Delta d_{L}}^{\mathrm{\Lambda-Sk}}$ of the inferred distance corrections: no clear bias is found either with respect to the sign of the correction or across redshift bins. In other words, while the mean corrections trace a coherent angular and radial structure, their associated uncertainties do not exhibit a systematic redshift-dependent trend, indicating that the signal is not driven by a redshift-dependent error bias but by the geometry of the reconstructed local expansion field.
@@
[…]
@@ 4 Conclusions
To constrain our model, we adopt a data–driven strategy based on the HAMLET CF4 reconstruction [112, 111] of the local density contrast and the full three–dimensional peculiar velocity field out to redshift $z\lesssim 0.15$ [110]. To suppress spurious small–scale nonlinearities, we smooth the original fields using a coarse–graining procedure over neighbouring voxels, ensuring that the resulting maps remain fully within the quasilinear regime. We then leverage the quasilocal formalism of the $\Lambda$–Szekeres solutions [101] to fit an approximate model to these smoothed data, and extract the quasilocal averages of both the expansion field and the spatial curvature in the local Universe (see Figs. 1 and 2).
@@
To assess the impact of these corrections on cosmological inference, we used the Pantheon+ SNe Ia subsample in the redshift range $0.023<z<0.15$ — the same range employed by the SH0ES collaboration to infer the Hubble constant. We performed a bootstrapping procedure over realisations of the employed CF4 reconstruction, and hence of the extracted quasilocal average expansion field, to determine for each supernova in the sample the directional corrections to the luminosity distances relative to the fiducial FLRW cosmology, along with their associated uncertainties. These corrections — which we find can modify supernova distances by up to 10% — were then incorporated into an MCMC analysis of a second-order cosmographic expansion’s parameter space to quantify the bias introduced by the local Universe expansion field on the estimation of $H_{0}$. We found that the inferred value of $H_{0}$ increases when considering the corrections predicted by our approximate $\Lambda$–Szekeres model, worsening rather than relieving the Hubble tension.
These findings align with a previous analysis [43], which suggests that local structures cannot reconcile early-Universe and late-Universe determinations of $H_{0}$. However, while [43] considered only a homogeneous, anisotropic model of the Laniakea supercluster, we account for the full network of filaments, clusters, and voids in the local Universe, providing a more complete assessment of how local structure affects supernova distances and, consequently, the inferred value of $H_{0}$. Nonetheless, we find that including the full complexity of the local cosmic web does not seem to alleviate the Hubble tension, indicating that the solution might lie beyond local-structure effects.
@@ Appendix A The quasilocal formalism for $\Lambda$-Szekeres models
The quasilocal formalism of $\Lambda$-Szekeres models, developed Sussman and Bolejko in [101] and expanded upon in [103, 102, 104], has been demonstrated to be particularly well adapted to describe multi-structured $\Lambda$-Szekeres models [103, 102].
Within this formalism, it is convenient to write the $\Lambda$-Szekeres metric in spherical coordinates as [101, 103, 102, 104]
@@
where $a=a(t,r)$, $\Gamma=\Gamma(t,r)=1+ra^{\prime}/a$ and $a^{\prime}=\partial_{r}a$. $\mathcal{K}_{q,\text{ini}}=\mathcal{K}_{q,\text{ini}}(r)$ is an a priori arbitrary function of $r$, whilst the Szekeres dipole function $\boldsymbol{W}(r,\theta,\phi)$ and its magnitude $\mathcal{W}(r)$ are given by
@@
where $X,Y$ and $Z$ are arbitrary functions of the radial variable. We can then define for the Hubble scalar, $H$, the curvature scalar, $\mathcal{K}$, and the density scalar, $\rho$, their respective “q-scalars” via [101, 103, 102, 104]
@@
Therefore, the q-scalars are proper-volume averages of their corresponding scalars with weight factor $\mathcal{F}$, which can be interpreted as a measure of the total gravitational binding energy within the volume [^7] This can be seen directly for spherically symmetric systems in the Newtonian limit, for which the identification is exact [49].. We can also define the exact fluctuations of any dynamical scalar $A$ as
@@
It can then be shown that the dynamics of the Szekeres solutions are now completely determined by the set of first-order evolution equations [101, 103, 102, 104]
@@
where we have defined $\Delta^{(\rho)}:=D^{(\rho)}/\rho_{q}$. The evolution equations (A.9)-(A.12) are completed by the following algebraic constraints on the dynamical variables [101, 103, 102, 104]
@@
Appendix B Coarse-grained structures in CF4
We apply a coarse–graining to the HAMLET CF4 fields [112, 111] in order to suppress small–scale nonlinear contamination and reconstruction artefacts that are known to arise in linear Eulerian frameworks (e.g., biased amplitude in the density contrast and locally unphysical values in voids). Concretely, we replace the field value at a given location in the original reconstruction grid with a sliding–window average over a cube of $16\times 16\times 16$ neighboring cells, i.e., over a top–hat kernel of linear size $\simeq 64~{\rm Mpc}/h$. This operation acts as an explicit low–pass filter: fluctuations on scales below $64~{\rm Mpc}/h$ are strongly attenuated, while the large–scale gradients (which dominate the quasilocal averages used in the $\Lambda$–Szekeres calibration) are retained. In practice, this brings the density contrast into the quasi–linear range required by our effective modelling, while smoothing the velocity field into a coherent streaming flow pattern consistent with the intended large–scale description.
Fig. 811 illustrate this directly by comparing the unsmoothed and coarse–grained maps of $\delta$, $v_{x}$, $v_{y}$, and $v_{z}$ (left vs right columns) across the three fundamental supergalactic planes. As expected, the coarse–graining washes out small–scale clumpiness and sharp cell–to–cell variations, most evident in the velocity components, while preserving the morphology of the dominant structures (extended overdense/underdense regions and the associated coherent streaming flows) across the full volume. In other words, the procedure reduces sensitivity to sub–voxel nonlinear motions and numerical artefacts, but does not erase the long–wavelength features that set the effective environment (and hence the quasilocal expansion/curvature fields) relevant for the distance–redshift corrections. This provides a visual validation that our smoothing is aggressive only below the target scale $\sim 64\,{\rm Mpc}/h$, and conservative with respect to the large–scale structure information that the $\Lambda$–Szekeres patchwork is designed to encode.
Figure 8: Density contrast field, $\delta$, as given in the CF4 Hamlet reconstruction (left column) and the corresponding result when employing the 16-cell coarse-graining procedure (right column). The top row shows the SGX-SGY plane, the middle row the SGY-SGZ plane and the bottom row the SGZ-SGY plane.
Figure 9: SGX component of the PV field, $v_{x}$, as given in the CF4 Hamlet reconstruction (left column) and the corresponding result when employing the 16-cell coarse-graining procedure (right column). The top row shows the SGX-SGY plane, the middle row the SGY-SGZ plane and the bottom row the SGZ-SGY plane.
Figure 10: SGY component of the PV field, $v_{y}$, as given in the CF4 Hamlet reconstruction (left column) and the corresponding result when employing the 16-cell coarse-graining procedure (right column). The top row shows the SGX-SGY plane, the middle row the SGY-SGZ plane and the bottom row the SGZ-SGY plane.
Figure 11: SGZ component of the PV field, $v_{z}$, as given in the CF4 Hamlet reconstruction (left column) and the corresponding result when employing the 16-cell coarse-graining procedure (right column). The top row shows the SGX-SGY plane, the middle row the SGY-SGZ plane and the bottom row the SGZ-SGY plane.
Appendix C Code routine validation
To validate our code for the $\Lambda$-Szekeres modelling, we employ a simple simulation of a toy model, the ‘Hubble bubble’, whose analytical behaviour is well understood. Specifically, we consider a spherical region of radius $r_{b}$ with a constant linear density contrast $\delta$, embedded in an otherwise flat FLRW background. The variation of the Hubble parameter inside the bubble is related to the density contrast and to the background expansion rate by [75]
$\frac{\delta H}{\bar{H}}=-\frac{\delta}{3}\Theta(\delta,\Omega_{m})f(\Omega_{m})\;,$ (C.1)
where $\bar{H}$ is the Hubble parameter outside the density contrast bubble, $\Theta\approx 1$ in the linear regime, and $f\approx\Omega_{m}^{0.55}$. To validate our modelling framework, we compare the distance–redshift predictions for light sources for (i) the effective $\Lambda$-Szekeres model, (ii) Newtonian perturbation theory, and (iii) the standard FLRW model. However, to build a Hubble diagram and compare the various models predictions, we need to specify the underlying distance–redshift relationship in the Hubble bubble model.